# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

### v1

@@ Introduction
As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as be able to consolidate the each graph of different node and edge types into a single question embedding. Initial approach was to concatenate of the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
@@
Following the initial exploration, we base our output layers largely on SHINE (Wang et al. 2021), where a dynamic graph of question nodes and cosine similarity edges is learned from the individual GCN outputs and the nodes are classified via label propagation in a final GCN layer. The training parameters are kept the same as provided by Wang et al.. An overview of the feature extraction, graph construction, and model is provided in Figure 1.
Experiments
Datasets
@@
This dataset was obtained from Najran University (Yahya et al. 2013), containing 596 questions with 3 class labels of different cognitive complexities.
@@
This dataset was released as part of the A12 Reasoning Challenge (Boratko et al. 2018), consisting 279 questions labeled with 3 different reasoning capabilities.
@@
This is a collection of 345 middle school science questions, labeled with 3 different type of answers required by the respective questions (Godea and Nielsen 2018).
@@
Bloom’s Taxonomy (Bloom and Krathwohl 1956) is one of the most well-established question classification frameworks used by educators. This dataset is a set of questions collected from different web sources, and manually classified into the six cognitive levels of Bloom’s Taxonomy (Yahya 2011).
@@
The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC (Hovy et al. 2001), about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set. These questions were manually labeled into 6 coarse and 50 fine class labels.
@@
Aside from TREC, for which the dataset split was obtained from HuggingFace ***https://huggingface.co/datasets/CogComp/trec, the train/test split was given by a pre-determined seed.
Performance Evaluation
@@
We compare the performance of our proposed model with selected baseline models in Table 2, including vanilla CNN and Bi-LSTM for non-graph classification models.
@@
PQ-GCN outperforms the baseline models in the macro-averaged evaluation on ARC, LREC, and TREC-coarse datasets, and second to the best performing model for NU and Bloom datasets. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed by Wang et al., showing the effectiveness of phrase-based features in improving question classification capabilities.
While performance on TREC-fine drops compared to its coarse label counterpart across all models is expected due to the large number of labels, our proposed model suffers from a larger difference compared to models such as HyperGAT or ME-GCN. TREC-fine was extensively labeled and was included in experiments to determine model capability in a more extreme and granular question classification task. In realistic educational settings, it is unlikely to have a question set that is labeled as intently as TREC-fine.
Considerations for System Implementation
In this section, we outline possible considerations for practical deployment of the trained model.
Compared with Transformer-based models (Vaswani et al. 2017) which scale quadratically with sequence length, GCNs offer a less computationally extensive alternative to text embedding-based classification. This allows for greater accessibility to educators without access to appropriate compute resources.
In addition, node classification-based question classification is quick to train as at every epoch, the corpus-level graph is passed through the network once to obtain the logits for all samples, unlike traditional neural networks where samples are passed into the network batch by batch. This allows for a quick refresh of the model based on any new data points or even a whole new dataset.
There is also no need for a lot of labels for label propagation via cosine similarity at the question-level dynamic graph—some may not have access to well-labeled set of questions nor the resources for manual labeling. This is a crucial factor for learning analytics systems, and the fact that the proposed model performs well in low-resource settings shows its potential in such environments.
The proposed features can be made modular to suit the question corpus in context i.e. each individual graph can be added or removed from the model input to build the dynamic question-level graph. This will allow for greater flexibility in deployment of the model in various settings.
Conclusion and Future Works
The proposed method on applying GCNs to question classification shows promising results in capturing the intricate word and phrase relationships within question text. The next steps involve refining the model for specific educational and AI-driven applications, such as adaptive learning platforms or automated tutoring systems. Future work will focus on optimizing the model for deployment at scale, ensuring they can handle large volumes of data and provide real-time classification on user-provided questions. In conclusion, we hope to provide sufficient basis for further research into feature extraction methods for graph neural networks, especially in automatic question classification.

### v2

@@ IV-B Model Architecture
As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding. Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
@@
Following the initial exploration, we base our output layers largely on SHINE [11]. An overview of model architecture is provided in Figure 2. Each graph is passed through their own 2-layer GCN. Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features. Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification. We provide the model and training parameters for the proposed PQ-GCN in Table II.
V Experiments
V-A Datasets
We first select a number of datasets used for question classification task to evaluate our proposed model on. The details are outlined in this subsection.
@@ V-A Datasets
This dataset was obtained from Najran University [28], containing 596 questions with 3 class labels of different cognitive complexities.
@@
This dataset was released as part of the A12 Reasoning Challenge [29], consisting 279 questions labeled with 3 different reasoning capabilities.
@@
This is a subset of 344 questions from the original dataset of science questions asked by teachers in real middle school classrooms, with a proposed taxonomy by expected answer types [7]. The subset was chosen by selecting 3 mutually exclusive class labels—’Very short answer’, ’Context sensitive’, and ’Answers will vary’.
@@
Bloom’s Taxonomy [30] is one of the most well-established question classification frameworks used by educators, based on different educational objectives. This dataset is a set of questions collected from different web sources and manually classified into the six cognitive levels of Bloom’s Taxonomy [^3] The dataset can be found at dx.doi.org/10.13140/RG.2.1.4932.3123.
@@
The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC [31], about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set. These questions were manually labeled into 6 class labels.
@@
Aside from TREC, for which the dataset split is obtained from HuggingFace  [^4] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request. Relevant statistics for the datasets are found in Table III.
V-B Performance Evaluation
TABLE VII: Keyword Analysis for Five Datasets
Frequency is calculated by % of questions in the label set that contains the keyword.
Dataset Label Top Occurring Keywords (Frequency)
NU Transfer would (11.8%), compare (6.4%), story (6.4%)
Knowledge words (12.5%), describe (9.5%), state (7.0%)
Application construct (16%), use (14%), would (12.0%)
ARC Linguistic matching likely (22.1%), best (16.0%), water (10.7%)
Basic facts best(16.7%), chemical (12.8%), new (10.3%)
Hypothetical experiment (30.0%), water (27.1%), students (20.0%)
LREC Very short answer energy (13.3%), mass (11.9%), density (11.2%)
Context sensitive explain (18.6%), think (12.1%), lab (11.3%)
Answers will vary energy (28.6%), explain (20.8%), would (13.0%)
Bloom Comprehension explain (37.8%), discuss (16.6%), describe (14.3%)
Knowledge list (22.4%), define (16.0%), five (12.8%)
Evaluation justify (23.6%), evaluate (19.2%), answer (10.4%)
Application calculate (14.6%), explain (13.6%), determine (11.1%)
Analysis differentiate (17.8%), compare (15.5%), analyze (12.5%)
Synthesis suggest (23.7%), propose (11.83%), design (10.8%)
TREC Entity name (9.8%), fear (4.9%), first (3.9%)
Description and abstract concept mean (5.23%), origin (4.2%), get (2.8%)
Human being name (11.6%), first (7.38%), president (5.8%)
Numeric value many (32.7%), year (8.2%), much (6.1%)
Location country (13.7%), city (12.0%), state (7.2%)
Abbreviation stand (48.4%), abbreviation (19.0%), mean (11.6%)
@@ V-B Performance Evaluation
We compare the performance of our proposed model against selected baseline models in Table VI, including vanilla CNN and Bi-LSTM for non-graph classification models.
The baseline models were selected by reviewing their contributions in graph neural networks for text classification [32], each using distinct feature engineering and graph construction methods for text data, as in Section II, as compared to works which have proposed architectural improvements for computational efficiency. With the exception of CNN and Bi-LSTM, for which parameters were set as outlined in Table IV and Table V, the rest of the baseline models have their model and training parameters set as provided by authors in the project repositories.
@@
PQ-GCN achieves the best macro-averaged F1 scores across most datasets, including NU (0.724), LREC (0.751), and TREC (0.801), while consistently delivering high precision and recall. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [11], showing the effectiveness of phrase-based features in improving question classification capabilities.
CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC. This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category. An analysis of most frequently occurring words have been included in Table VII, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords. For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest”, “propose”, and “design” that does not appear in questions of other classes. While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would”, “water”, and “energy”.
TensorGCN performs with the highest F1 score on TREC. TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets. While this brings about a 0.004 point gain in F1 score over PQ-GCN, it also results in a longer graph construction time with increasing dataset size due to increased training time of the LSTM model, as compared to the proposed methodology of using pre-trained embedding models.
Bi-LSTM has a consistently average performance across the datasets, which is expected of a simple sequence model. TextGCN shows apt results for NU and ARC datasets where its F1 scores are second only to PQ-GCN, but there is room for improvement in leveraging text features other than PMI and TF-IDF when constructing the graph, especially when it comes to text with niche characteristics such as questions. Text-Level GNN shows inconsistent poor performance across the datasets, as its learnable word-word edge weights are highly dependent on relative proximity of keywords and the ability of the message passing mechanism to disambiguate polysemous words based on word proximity. HyperGAT also shows inconsistent performance, possibly attributed to its word-word edge features, which are extracted by segmenting a text at sentence level and treating each sentence as a “window”, as opposed to a sliding window in PMI calculation. Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on. ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQGCN outperforms both models in F1 scores across all datasets.
VI Conclusion
We develop and evaluate PQ-GCN, adapting GCNs to question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions. The proposed method shows promising classification performance compared to baseline models, capturing the intricate word and phrase relationships within question text. Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning, augmented with artificial intelligence. Finding suitable feature extraction methods is crucial in tackling this task, and we hope to provide sufficient basis for further research into innovative feature extraction methods for automatic text classification with graph neural networks.